OpenStudy (anonymous):

How to graph R(x) = 2(x+1)(x-5)/(x+3)(x-1) What are the x and y intercepts along with the hole?

OpenStudy (shamil98):

x-intercepts , set r(x) = 0 and solve for x. y-intercepts, after you find the x-intercepts, just plug them back into the function to find the corresponding y-intercepts. To graph it, make a table, r(x) = y value so you just input x values and find the corresponding y-values and then graph it..

OpenStudy (anonymous):

so X intercepts are (3,0), (5,0) and y intercept (0,-10)?

OpenStudy (anonymous):

I could really use your help, I have a final tomorrow.

OpenStudy (shamil98):

No those are wrong.

OpenStudy (shamil98):

One of them is right though, x- intercept --> 5 and its corresponding y intercept is 0 . so (5,0)

OpenStudy (anonymous):

Then what would the other one be? The y intercept. Would the graph look like to two curved lines diagonal from one another?

OpenStudy (shamil98):

The other x-intercept is (-1,0)

OpenStudy (shamil98):

To graph it, make a table, and then plot the x , y values and connect them.

OpenStudy (anonymous):

So using the asymptotes as a guide, right?

OpenStudy (anonymous):

I know the horizontal asymptote would be 2, right? What would the vertical be?

OpenStudy (shamil98):

Vertical asymptotes would be found by setting your denominator equal to 0 and solving for x.

OpenStudy (shamil98):

(x+3)(x-1) = 0 x = ? , x= ?

OpenStudy (anonymous):

so -3 and -1?

OpenStudy (anonymous):

We are almost done ;)

OpenStudy (shamil98):

-3 , and postive 1. x - 1 =0 x = 1

OpenStudy (anonymous):

Can I ask one other quick question?

OpenStudy (shamil98):

Go for it.

OpenStudy (anonymous):

What is the probability that a person is in favor of a school policy change given that she is a student? Total amount of students = 619 Those in favor of Students - 353 Faculty - 34 What is the probability?

OpenStudy (shamil98):

619 + 34 = 653 says a person , so add students + faculty to get the max number of people than divide the total number by those in favor. 353/653

OpenStudy (anonymous):

Thank you so much.

OpenStudy (shamil98):

you're welcome.